What is the exact value of sin 2theta given sin theta = (5/13)?
Answer by stanbon(75887)

What is the exact value of sin(2theta) given sin(theta) = (5/13)?
Note: cos(theta) = sqrt[13^2-5^2]/13 = sqrt(144)/13 = 12/13
------
Your Problem::
sin(2t) = 2*sin(t)cos(t) = 2(5/13)(12/13) = 120/169
===============
Cheers,
Stan H.
---------------